OpenStudy (anonymous):

Write 1-2log7^x as a single logarithm

OpenStudy (ranga):

2log(7^x) = log(7^(2x)) 1 can be written as log(10) 1-2log7^x = log(10) - log(7^(2x)) Then you can combine the logs using the identity: log(A) - log(B) = log(A/B)